OpenStudy (anonymous):

Locate the discontinuities of the function. y = ln((tan(x))^2)

OpenStudy (anonymous):

\[\ln ((\tan (x))^2)\]

OpenStudy (anonymous):

this is what I got in Maple... I hope it will help

OpenStudy (nikvist):

Function has discontinuities, when: \[\tan{x}=0\] or\[\tan{x}\rightarrow\pm\infty\] \[\Rightarrow\quad x=k\frac{\pi}{2}\enspace,\enspace k\in\mathbb{Z}\]